Probability that an office chosen randomly has two doors and three windows
1
Expert's answer
2021-07-22T05:32:48-0400

"\begin{matrix}
 & Number \ of\ doors\\
 Number\ of\ windows & \def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & 1 & 2 & 3 \\ \hline
 1 & 5 & 5 & 3 \\
 \hdashline
 2 & 1 & 1 & 2 \\
 \hdashline
 3 & 3 & 0 & 0 \\
 \hdashline
 4 & 7 & 3 & 0 \\
 \hdashline
\end{array}
\end{matrix}"


Total number of offices is


"5+5+3+1+1+2+3+0+0+7+3+0=30"

Number of offices having two doors and three windows is .


Then the probability that an office chosen randomly has two doors and three windows is equal to zero.